Which of the following describes the effect on the area of a triangle when the height is doubled and the base remains the same.
The area is reduced by a factor of 1/2.
The area is doubled.
The area remains constant.
The area is increased by a factor of 4.

Answer by Theo(13342)

i believe the area will be doubled.

for example:

assume a base of 10 and a hight of 6

area = 1/2 * b * h = 1/2 * 10 * 6 = 30.

when you double the height, the area becomes 1/2 * 10 * 12 = 60.

if you just choose alphas rather than numbers, you get:

A = 1/2 * b * h.

then the height is doubled, it becomes 2 * h, so the formula becomes A = 1/2 * b * 2 * h which results in A = b * h which is double 1/2 * b * h.